(1 point) a. Find the most general real-valued solution to the linear system of differential equations t = 1 -6 x1 (t) = C1I AM STRUGGLING HEAVILY ON THIS PROBLEM PLEASE HELP ME WITH THE CORRECT ANSWERS AND NEAT WORK PLEASE PLEASE AND THANK YOU
Answer #1

Ques 2!= 5* -;]? A= [] First, we will find eigenvalue of A:- -bord -4 -6-1 -6 tet land +40 =0 Sol d=-12 + |144-160 2 d= - 12o بالا - 22. . Cosatt isinat Now, Rq » Rq -R, [6.13] So, sataisa So, 24 = 36, = -2ixa 112=362 So, the general solution is X=50) we can write two linerely indebendent solution- X(t)= e-bt C, 1-25142A Cosat tal glosat singt (b) Here the eigin values aS9 it is a spiral Х), 6 bе (у г<ә, 72, 73<чЯ (44)